On a 235 mile trip, Clara drove 25 miles before she could set the cruise control to 60 miles per hour. How long did she drive with cruise control?
Found 2 solutions by psbhowmick, fractalier:
Answer by psbhowmick(878)

235 - 25 = 210 miles were yet to be driven when the cruise control was engaged.
So Clara drove with cruise control for 210/60 = 3.5 hours

Answer by fractalier(6550)

Well if she drove 25 miles, she had 235-25 = 210 miles left.
Since distance = rate times time, or d = rt, we have
t = d/r = 210/60 = 3.5 hours on cruise control